MODS1 Commissioning Run 5: UTC 2011 February 17

Observer(s): R. Pogge & J. Antognini (OSU), M. Pedani (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Little (LBTO)
Instrument Support: J. Morris (LBTO)
Team MODS Support: J. Mason (OSU)

Summary

Bad weather all night - high winds and high humidity above limits. A total loss for on-sky time. Used the closed-dome time to work through calibrations, test some code enhancements, and get mileage on the instrument and scripting engine. Re-measured and implemented the IMCS collimator TTF zero points at Zenith.

Got our first crack at tuning up flat-field and comparison-lamp calibration parameters for an MOS mask. Prototype scripts tested and ready when we get on sky next.

Preparations

Inspecting the mask system during the day with some suggestions from Tom O'Brien. I pulled the mask that was in the problem position last night and can find no obvious flaws. Since the Stone I mask is the wrong scale, we remounted the Stone F mask in another frame and installed it in the position occupied last night by the Stone I mask. The elevation/rotator position test will be repeated since later tonight may be bad weather if the forecast is to be believed.

MODS computers moved from the lab mini-rack into their new home in the full-size rack in the Level 2 Computer Room B. Jerry Mason has come up with a workable configuration, with remote assistance from Kyle Foster.

Installing a copy of the pyraf script to do MOS mask alignment on the obsN machine (for user MODSeng at first). Stephen Hooper successfully installed pyraf and numpy on the obsN cluster. Thanks!

MODS filled at 4:30, still not running out in either dewar in 24h

Details

Afternoon - Mason moving the MODS computers into the LBT rack in Level 2 Computer Room B. This includes installation of an 8-port Raritan KVMoIP to replace the two smaller single-channel units connected to the ICs in the lab. Some reconfiguration of the rack was required to accommodate the power strips and cable management rails with the deeper "draft" of the Dell server ("mods1") and the fold-out monitor and keyboard set. Work started early (MODS was shutdown after the end of the night in preparation for this).

Took out the Stone F field mask and inspected it - nothing obviously wrong. However, I remounted it in another frame, then extracted the Stone I mask which has scale problems, and put the Stone F mask in its new frame into its place in Position 20.

All times are UTC unless otherwise noted.

0001 - Wind still strong, 14m/s form the SSW with gusts to 18m/s, Temp 1°C, humidity 48% and rising all day. Clouds starting to come in from the north and form high overhead. The forecast is not looking promising. We might get a start if the wind cooperates.

0100 - Winds gusting over 20m/s, max so far 23m/s. No chance of opening right away. Clouds increasing but still lots of blue shinging through. Maybe sunset will be pretty...

0115 - ... or not, the clouds to the WNW are very thick and there wasn't much color. Oh well.

0200 - Winds still high, though starting to decline a little from the past hour. Humidity has risen to 66%, temperature has fallen to a little below 0°C. Clouds now coming in more heavily from the NNW

Not looking hopeful. A night for reducing data, testing slit-mask sensors, and other diversions.

0430 - Windy, up to 22m/s SSW, humidity 85%, temp -1°C. Still just clouds blowing by overhead

IMCS Collimator TTF Zero-Points

0530 - The instrument configuration file contains nominal collimator TTF values that should put the IMCS laser spot on or very near the quad-cell center when the telescope is at Zenith in rotator PA=0° (LDG rotAngle=210°). When changing between configurations, this should give the IMCS a head-start on locking on.

To measure these requires switching among the 9 instrument modes : {dual,blue-only,red-only}x{imaging,grating,prism} and then requesting IMCS lock-on. After lock, a short 8x8 binned 1Kx1K bias image is acquired, storing the collimator TTF actuator values (COLTTFA/B/C), the date, time, telescope configuration, instrument air and truss temperatures, etc. as the overall record of the measurement.

These get stored on the mods1data computer under Calibrations/TTFs/ in a directory named CCYYMMDD for the date of the measurement.

The TTF values get put onto the active MODS1 configuration tables on host mods1: Config/MODS1/bconfig.tab and rconfig.tab, copies of which should be stored with the images and summary data analysis on mods1data

The measurment is made from the mods1data machine using the the imcsTTFs.pro script. Execution time is about 12 minutes, excluding initial setup of the telescope and instrument. It can be done with the enclosure lights on, but the instrument should be in its dark mode (hatch closed, calibration tower in the beam, and all external hatches closed).

The measurements tonight are as follows:
   Channel  Mode    Disp       A      B      C
   ---------------------------------------------
    Blue    Direct  Imaging  16997  17958  20150
                    Grating  16963  18074  20069
                    Prism    16934  17948  20221
            Dual    Imaging  17165  17876  20065
                    Grating  17149  17980  19975
                    Prism    17092  17876  20136
   ---------------------------------------------
    Red     Direct  Imaging  15891  14877  12617
                    Grating  15915  14954  12517
                    Prism    15844  14979  12563
            Dual    Imaging  15828  14919  12639
                    Grating  15858  14975  12552
                    Prism    15795  15011  12580
   ---------------------------------------------

Spectral Calibrations - Flat Fields & Comparison Lamps

Taking spectral flat fields with the new calibration scripts. ...

0715 - Weather update: 100% humidity, wind 18m/s gusts to +20m/s out of the SSW, Temperature -2°C. Clouds still above us but thick enough there is no moon

Long-Slit Spectral Flats

Taking grating flats with the LS5x60x0.6 long slit and the grflats.cal script. This is the same script created in January, but now with the hooks for the beta-release script engine.

A minor glitch with the data-transfer system at the start, a residue of some maintenance work on the ICs related to the rack transfer cleanup that we forgot to clear (normal procedure would be to reboot, but we forgot - the corrective was to reboot).

MOS Mask Spectral Flats

MOS masks often have wider slits especially if working on faint objects to the outer edges of the MODS field-of-view where the images begin to get aberrated. They also have alignment star holes which admit more light in a spectral pixel than a slit. This reduces the saturation threshold for the combination of lamp choice and exposure time.

Worked out a basic mosflats.cal script in dual mode for the Abell1689 commissioning mask. This mask has 1-arcsec wide slits and 4x4 arcsec alignment star holes. The latter saturated, sometimes to the bleed/streak level, at the normal long-slit mask lamp/exptime parameters. Tweaked up the exposure times.

The best combination seems to be
  • Blue: 1 QTH lamp, sequences with ND1.5 and UG5 filters
  • Red: VFLAT at level 10 and Clear Filter, but about half the exptime of the long-slit flats

There are no good lamp combos that can be shared by both channels, requiring serialization of the data acquisition.

0930 - Weather update: the cloud layer has descended to our level, we're just in the base of the deck. Winds and humidity are over their limits, temperatures are -3°C

MOS Comparison Lamps

This is a variation on the long-slit lamps. The test script is gmoslamps.cal Slight alterations in exposure times were needed for the larger apertures as with the flats. However, unlike the flats there are a couple of good combos that let us take parallel lamp spectra.

Used the Abell1689 commissioning mask for this development work.

Prism comparison lamps were taken using the prototype script pmoslamps.cal, using the prlamps.cal script as a template. Some minor exposure time tweaks were required for the red, but otherwise the exptimes are similar. Again there are fruitful combos that permit concurrent calibration spectra in both channels.

MOS Spectral Flats

Using the prflats.cal as a template. One "discovery" is that the MOS prism mode really wants a 4Kx3K subframe ROI for readout. Implemented and tested, gives enough room for all spectra out to the edges of the mask.

Prototype script is pmosflats.cal

Repeated pmoslamps.cal and pmosflats.cal with the 4Kx3K ROI. Much better.

-- RichardPogge - 16 Feb 2011
Topic revision: r7 - 17 Feb 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback